Please state the coordinates of the vertex.
12y+x^2-4x=-16

Answer by lwsshak3(11628)

Please state the coordinates of the vertex.
12y+x^2-4x=-16
x^2-4x=-12y-16
complete the square:
(x^2-4x+4)=-12y-16+4
(x-2)^2=-12y-12
(x-2)^2=-12(y+1)
This is an equation of a parabola that opens down.
Its standard form: (x-h)^2=-4p(y-k)^2, (h,k)=(x,y) coordinates of the vertex
For given equation:12y+x^2-4x=-16
vertex: (2,-1)